OpenStudy (anonymous):

To which subsets of the real numbers does the number √42 belong? A. whole numbers, natural numbers, integers B. irrational numbers C. rational numbers D. whole numbers, integers, rational numbers

OpenStudy (anonymous):

@goformit100 @hkm91

OpenStudy (goformit100):

@RoxxisaUNICORN What answer do you thing is correct ?

OpenStudy (anonymous):

idk (/./) eith a or d?

OpenStudy (anonymous):

@Notamathgenius

OpenStudy (goformit100):

yes you are correct

OpenStudy (anonymous):

Lol I said a OR d :p

OpenStudy (anonymous):

b irrational

OpenStudy (anonymous):

a rational number is a number that can be expressed as a simple fraction e.g. 4=8/2 or 5.5=11/2 in your case the number cannot be expressed as a simple fraction (i.e using whole numbers) you can't get rid of the square root so it is irrational only. a natural number is the positive whole numbers , integers are the positive and negative whole numbers.

OpenStudy (anonymous):

might any one know this her ?? is something like mine hope u dont mind